OpenStudy (anonymous):

please help the asignment is due in 1 minute please help asap will fan and medal

OpenStudy (anonymous):

Larry’s Landscaping charges $235 for spring cleanups and $30 for weekly lawn maintenance. Joe’s Landscaping charges $185 for spring cleanups and $35 for weekly lawn maintenance. The system that models this situation is given, where c is the cost of lawn maintenance and w is the number of weeks. c = 235 + 30w c = 185 + 35w The solution to the system is (10, 535). Which interpretation correctly describes the solution to the system of equations? A. Larry’s Landscaping will charge more money on the tenth week by charging $535. B. Joe’s Landscaping will charge more money on the tenth week by charging $535. C. The cost for lawn maintenance is the same, $535, for both landscaping companies after 10 wk. D. A customer can have his or her lawn maintained for a maximum of 10 wk. The customer will pay a total of $535.

OpenStudy (anonymous):

@angelic._. @AnswerMyQuestions @Blu-Girl @HaileyD